# Learning to Route LLMs with Confidence Tokens

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive performance on several tasks and are increasingly deployed in real-world applications. However, especially in high-stakes settings, it becomes vital to know when the output of an LLM may be unreliable. Depending on whether an answer is trustworthy, a system can then choose to route the question to another expert, or otherwise fall back on a safe default behavior. In this work, we study the extent to which LLMs can reliably indicate confidence in their answers, and how this notion of confidence can translate into downstream accuracy gains. We propose Self-REF, a lightweight training strategy to teach LLMs to express confidence in whether their answers are correct in a reliable manner. Self-REF introduces confidence tokens into the LLM, from which a confidence score can be extracted. Compared to conventional approaches such as verbalizing confidence and examining token probabilities, we demonstrate empirically that confidence tokens show significant improvements in downstream routing and rejection learning tasks.
11footnotetext: Work done during the internship at Apple.
@@ 1 Introduction
Recent years have seen an explosive growth in the deployment of Language Models (LLMs), in forms ranging from helpful conversational agents [1, 49, 56], to frameworks for automating software workflows [15, 53], to tools tailored to domain-specific tasks. LLMs have been used to summarize doctor-patient interactions [29, 47, 3], answer questions from students [45, 54], and provide customer support [28]. As LLMs are given more agency in settings of increasing consequence, it becomes crucial to know when an output is reliable [55]. Given knowledge of the trustworthiness of an output, systems can proactively prevent faulty behavior by seeking a second opinion (e.g. by routing to a more costly LLM), or choosing to abstain from answering (e.g. instead defaulting to a safe behavior).
@@
However, the state-of-the-art in LLMs face challenges with providing an accurate estimate of confidence in whether a prediction is correct. Several works identify that token probabilities (derived from a softmax over logits) are not well-aligned, or calibrated, with the actual probabilities of correctness [22, 12, 52]. Others have found that modifying the prompt by asking the LLM to verbalize its confidence yields slightly better results [48, 50, 36]; however, these results may be subject to the dataset and prompt engineering, often leading to unstable or unreliable results. Since LLMs are typically trained using a cross-entropy loss, they can overfit on accuracy rather than calibration, often leading to overconfidence and misalignment with real-world distributions. Thus, a natural question arises: how can we train LLMs to accurately estimate their own confidence levels, while preserving their performance on tasks of interest?
@@
Figure 1: The routing pipeline of the Self-REF framework based on end-to-end confidence prediction.
@@
To answer this question, we introduce Self-Reflection (Self-REF), a lightweight framework for fine-tuning LLMs to accurately assess confidence in their predictions while preserving their performance on relevant downstream tasks. Self-REF integrates naturally with any LLM backbone, allowing the model to condition its assessment of confidence on its generated outputs. It consists of three steps: (i) Confidence token annotation: the training dataset is generated by labeling each instance with confidence tokens based on the correctness of the base LLM. When the base LLM provides correct responses to the input text, we augment the instances with confident (<CN>) tokens. Conversely, if it responds incorrectly, we augment the instances with unconfident (<UN>) tokens. (ii) Self-REF fine-tuning: The base LLM is fine-tuned on the augmented training data, so that one would expect a confidence token to be generated following its predicted output. (iii) Confidence score extraction: To obtain a continuous confidence score, one computes the probability of the <CN>token, normalized by the sum of the probabilities of the <UN>and <CN>tokens.
@@ 2 Related Work
In this section, we briefly introduce related studies on uncertainty quantification, LLM routing, and LLM rejection learning. More discussion is provided in Appendix B.
@@ 2.1 Uncertainty Quantification in LLMs
Recent studies in LLM uncertainty have explored leveraging internal features like logits and attention scores to estimate uncertainty without additional training [61, 36, 50]. Post-processing methods such as temperature scaling [14], prompting LLMs to verbalize confidence scores [48], and instruction learning/pretraining [60, 42, 32, 35, 10] have also shown improvements in calibration. In contrast to these methods, Self-REF allows the LLM to learn to estimate its confidence at fine-tuning time. Furthermore, this internal notion of confidence is aligned against correctness rather than the consistency of responses after re-sampling, as is studied in [57, 37, 34]. As noted in [22, 51], uncertainty scores derived from LLMs may exhibit a weak correlation with actual prediction correctness, and the most consistent answers are often not the most correct. Finally, while much of the prior work focuses on calibration, our work goes beyond it and focuses on the correlation between downstream utility of model and confidence scores for LLM routing and rejection learning.
@@ 2.2 LLM Routing
In confidence-based query-routing scenarios, one line of research trains additional classifiers to route queries to language models (LLMs) based on observed performance metrics and routing data, forming routing pipelines with trained multi-agent systems or extra routers to boost performance [13, 43, 46, 24, 59]. Another approach leverages LLM cascades, invoking increasing complex LLMs depending on trained scoring functions which predict whether an answer is correct, and consistency of responses upon resampling [8, 58]. Unlike prior work in LLM cascades, Self-REF trains the estimate of the probability of correctness into the LLM itself instead of an external model. Additionally, the estimates of confidence from Self-REF are aligned against the correctness of the predictions, rather than the consistency. Furthermore, Self-REF computes the confidence in the autoregressive decoding step, allowing for LLM itself to condition on its generated output. This dynamic approach allows for more accurate and adaptive routing decisions based on real-time confidence assessments.
@@ 2.3 LLM Rejection Learning
Learning to reject [11] was originally introduced to equip the Bayes classification model with a rejection option for predictions. Despite advancements in LLMs, their predictions can still be prone to uncertainty due to inherent knowledge limitations. This highlights the need for LLMs to have the ability to refuse to answer, especially in situations where it is infeasible to fallback upon more advanced models. LLMs can be trained for the rejection task through rejection knowledge injection [7], an additional rejection LLM agent [38, 39], and rejection-oriented loss adaptations [33, 40]. However, such approaches involving extra knowledge injection or loss adaptation may degrade the base performance of the LLM. In this work, we aim to equip LLMs with the capability for rejection learning based on their confidence levels, without the need for designing new loss functions, thereby preventing any potential degradation in the performance of downstream tasks.
@@ 3.2 Self-REF
The Self-REF framework fine-tunes a base model $\mathcal{M}{}$ to use confidence tokens and also learn their embeddings. The framework involves (i) confidence token annotation, where data are augmented with confidence tokens, and (ii) Self-REF fine-tuning on the augmented data, and (iii) confidence score extraction.
@@
1:Base LLM $\mathcal{M}{}(\cdot)$ and training data
2:$\mathcal{D}_{\mathrm{train}}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N_{\text{train}}}$.
3:Augmented data $\mathcal{D}^{\prime}_{\mathrm{train}}$ with unconfident samples and the confident samples. Use $\mathcal{M}{}(\cdot)$ to make predictions $\widehat{y}^{(i)}$ for the given input query $x^{(i)}$.
4:Create a set of unconfident samples
5:$\mathcal{D}^{\prime}_{\mathrm{train},\texttt{<UN>}}=\{(x^{(i)},\widehat{y}^{(i)}\texttt{<UN>}):y^{(i)}\neq\widehat{y}^{(i)}\}_{i=1}^{N_{\text{train}}}$.
6:Create a set of confident samples
7:$\mathcal{D}^{\prime}_{\mathrm{train},\texttt{<CN>}}=\{(x^{(i)},\widehat{y}^{(i)}\texttt{<CN>}):y^{(i)}=\widehat{y}^{(i)}\}_{i=1}^{N_{\text{train}}}$.
8:Mix the unconfident and confident data to form the combined augmented dataset, subsampling the unconfident samples with tunable proportion $\alpha\in[0,1]$:
@@
Note that the unconfident and confident tokens are generated end-to-end conditioned on both the prefix query and the answer that the LLM itself generated, in contrast to prior work [43, 13, 46] which typically routes based on the query alone with an extra router. By fine-tuning with the LLM’s own predictions on the training set, Self-REF customizes the notion of uncertainty to the LLM itself, rather relying on the inherent uncertainty of a query.
@@
After training, one can extract a continuous confidence score $c_{\mathcal{M}}(x^{(i)},\\ \widehat{y}^{(i)})\in[0,1]$ by getting the token probability of <CN>, normalized over the sum of the <UN> and <CN> token probabilities: $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})=\frac{P(\texttt{<CN>})}{P(\texttt{<UN>})+P(\texttt{<CN>})}$. This continuous score gives us some control over navigating different tradeoffs in downstream settings.
@@ 3.3 Learning to Route and Reject
In confidence-based routing, answers where the smaller fine-tuned LLM is uncertain, i.e., $c_{\mathcal{M}}(x^{(i)},\\ \widehat{y}^{(i)})<t$ for some chosen threshold $t\in[0,1]$, will be routed to a more powerful LLM $\mathcal{M}_{\text{Large}}(\cdot)$. For answers with $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\geq t$ (i.e., certain), the answers from the local LLM $\mathcal{M}{}(\cdot)$ will be directly returned to the user.
@@
To study this question, we create an evaluation set where half of the samples do not contain the ground truth, i.e., we remove all ground-truth information from $x^{(i)}$, and replace the label with “none of the above,” $y^{(i)}=\emptyset$. Ideally, the model’s confidence in any of the remaining answers should be low, and we can evaluate how good a policy of choosing to abstain from answering the question does against this dataset. Thus, if $\mathcal{M}{}(\cdot)$ has a low confidence, i.e., $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$ for a chosen threshold $t\in[0,1]$, we treat it as if the model were to abstain from answering, i.e., $\widehat{y}^{(i)}=\emptyset$.
@@ 4.1 Datasets and Baselines
Dataset
All experiments are conducted on four public datasets: Massive Multitask Language Understanding (MMLU) [17, 18], OpenbookQA [9], GSM8K [9], and MedQA [25] datasets. The MMLU dataset consists of multiple-choice questions covering a wide range of knowledge domains with 57 distinct tasks. The OpenBookQA is a question-answering dataset designed to test deep understanding through multi-step reasoning, common knowledge, and text comprehension, similar to open-book exams. The GSM8K is a dataset containing graduate school math questions, where each question is collected from the Math World Problem Repository [44]. The MedQA is a multiple-choice open domain question answering dataset for solving medical problems collected from professional medical board exams. More details can be found in Appendix C.
Baselines
We compare the Self-REF framework to four state-of-the-art baselines for measuring model confidence, following the settings from existing work. Details about the prompts and hyper-parameters utilized in these baselines are given in Appendix D. We list the baselines as follows:
  • Verbalizing Uncertainty [48]: Prompt the LLM with in-context learning to yield confidence scores between 0 and 1. This encourages the LLM to generate its response while simultaneously estimating the confidence level of the response.
  • Verbalizing Yes/No Tokens [48]: Prompt the LLM with a question of whether the model is confident in its answers. The confidence score is calculated by normalizing the probabilities of the “Yes” and “No” tokens by their sum.
  • Zero-shot Logits [36]: Extract the probability of the generated prediction from the LLM without fine-tuning. In free-text generation questions, we average the probabilities of the tokens from the output answers. In the closed-form generation questions with single-token answers (e.g., multiple-choice questions), we directly select the probability of the predicted token.
  • Fine-tuned Logits [35]: Extract the probability of the generated prediction from a fine-tuned LLM. All methods for computing the probabilities of the outputs are identical to the “Zero-shot Logits” baseline. The confidence score is determined by the probabilities.
4.2 Experimental Settings
Confidence-based Routing Task
In the routing task, we aim to accurately route instances that the smaller LLMs (Llama3-8B-Instruct and Mistral-7B-Instruct) struggle with to the larger LLMs (Llama3-70B-Instruct) for improved performance. All experiments utilize Llama3-70B-Instruct with only its strong in-context learning capabilities during instance routing. The routing decisions are determined by the probabilities of the predicted confidence token <CN> and <UN> generated from Self-REF. To characterize the trade-off between cost and performance, we analyze performance along several possible routing thresholds $t$. In particular, for a set of confidence scores $\mathcal{C}_{\mathcal{M},\mathcal{D}}=\{c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\}_{i=1}^{N}$, we set $t$ at quantiles, where $t=\mathcal{Q}_{p}\left(\mathcal{C}_{\mathcal{M},\mathcal{D}}\right)$ for $p\in\{0,1,2,\dots,20\}$. Instances are routed to the larger LLM when $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$, so these thresholds correspond to routing approximately 0%, 5%, 10%, …, 100% of total queries to the larger LLM.
Confidence-based Rejection Learning Task
Implementation Details
We select Llama-3-8B-Instruct [49] and Mistral-7B-v0.3-Instruct [23] models as the pre-trained local base LLMs for Self-REF. Llama-3 70B model [49] is the larger, more powerful LLM to which unconfident queries are routed. Parameter-efficient fine-tuning is performed using LoRA adapters with ranks 2, 4, and 8 following the observations from [16]. Under at most eight epochs of training, all other hyperparameters on model training are decided by grid search with the perplexity loss on validation sets. More details of hyper-parameters are in Appendix D.
5 Experiment Results
5.1 Routing Performance (RQ1)
@@
Figure 2: Overall accuracy vs. routing rate in confidence-based routing task. Routing is from local LLMs Llama3-8B-Instruct (row 1) and Mistral-7B-Instruct (row 2) to Llama3-70B-Instruct on MMLU (column 1), OpenbookQA (column 2), GSM8K (column 3), and MedQA (column 4). Routing rate of 0 gives pure local LLM performance, and routing rate of 1 gives pure Llama3-70B-Instruct performance. Self-REF consistently outperforms baselines.
Overall Accuracy
Confidence-based routing using Self-REF consistently achieves the best accuracy vs. routing rate trade-off for all four datasets (MMLU, OpenbookQA, GSM8K, MedQA) and both local LLMs (Llama3-8B-Instruct model and Mistral-7B-Instruct) (Figure 2). Using Self-REF on Llama3-8B-Instruct for MMLU, confidence-based routing of just 39% of queries can achieve comparable performance to routing to the Llama3-70B-Instruct model alone. Similar observations hold for the OpenbookQA, GSM8K, and MedQA datasets, where routing the Llama3-8B-Instruct model with rates 49%, 65%, and 40%, respectively, can yield comparable performance to Llama3-70B-Instruct alone. For Self-REF on Mistral-7B-Instruct, the minimum routing rates to achieve parity with Llama3-70B-Instruct are 70%, 50%, 75%, and 70% for MMLU, OpenBookQA, GSM8K, and MedQA datasets, respectively.
Efficiency Gains
Smaller local models often have lower latency (sec) during inference as well as lower monetary cost. Thus, routing only a fraction of queries to larger, more powerful models can significantly improve the efficiency of the overall system compared to relying entirely on the more powerful models. For each small local LLM (Llama3-8B-Instruct and Mistral-7B-Instruct), we select the best routing rate as the minimum routing rate that achieves comparable performance to Llama3-70B-Instruct (see Figure 2 for the trade-off curves). Without sacrificing performance compared to Llama3-70B-Instruct, we observe that Self-REF achieves a 2.03×2.03$\times$2.03 × latency improvement for Llama3-8B-Instruct on MMLU and a 2.00×2.$0$ latency improvement for Mistral-7B-Instruct on OpenbookQA datasets (as shown in Table 1).

### v3

@@ Abstract
Large language models (LLMs) have demonstrated impressive performance on several tasks and are increasingly deployed in real-world applications. However, especially in high-stakes settings, it becomes vital to know when the output of an LLM may be unreliable. Depending on whether an answer is trustworthy, a system can then choose to route the question to another expert, or otherwise fall back on a safe default behavior. In this work, we study the extent to which LLMs can reliably indicate confidence in their answers, and how this notion of confidence can translate into downstream accuracy gains. We propose Self-Reflection with Error-based Feedback (Self-REF), a lightweight training strategy to teach LLMs to express confidence in whether their answers are correct in a reliable manner. Self-REF introduces confidence tokens into the LLM, from which a confidence score can be extracted. Compared to conventional approaches such as verbalizing confidence and examining token probabilities, we demonstrate empirically that confidence tokens show significant improvements in downstream routing and rejection learning tasks.
Machine Learning, ICML
@@ 1 Introduction
Recent years have seen an explosive growth in the deployment of large language models, in forms ranging from helpful conversational agents (Achiam et al., 2023; Touvron et al., 2023; Yang et al., 2024), to frameworks for automating software workflows (Harrison, 2022; Wu et al., 2023), to tools tailored to domain-specific tasks. LLMs have been used to summarize doctor-patient interactions (Krishna et al., 2021; Jeong et al., 2024), teach students (Schmucker et al., 2023; Xiao et al., 2023), and provide customer support (Kolasani, 2023). As LLMs are given more agency in settings of increasing consequence, it becomes crucial to know when an output is reliable. Given knowledge of the trustworthiness of an output, systems can proactively prevent faulty behavior by seeking a second opinion (e.g. by routing to a more costly LLM), or choosing to abstain from answering (e.g. instead defaulting to a safe behavior).
@@
However, state-of-the-art LLMs face challenges with providing an accurate estimate of confidence in whether a prediction is correct. Several works identify that token probabilities (derived from a softmax over logits) are not well-aligned, or calibrated, with the actual probabilities of correctness (Huang et al., 2023; Detommaso et al., 2024; Wightman et al., 2023). Others have found that modifying the prompt by asking the LLM to verbalize its confidence yields slightly better results (Tian et al., 2023; Turpin et al., 2024; Mahaut et al., 2024); however, these results may be subject to the dataset and prompt engineering, often leading to unstable or unreliable results. Since LLMs are typically trained using a cross-entropy loss, they can overfit on accuracy rather than calibration, often leading to overconfidence and misalignment with real-world distributions. Thus, a natural question arises: how can we train LLMs to accurately estimate their own confidence levels, while preserving their performance on tasks of interest?
@@
Figure 1: Self-REF attaches a confidence token to each prediction. Routing proceeds based on the confidence.
@@
To answer this question, we introduce Self-Reflection with Error-based Feedback (Self-REF), a lightweight framework for fine-tuning LLMs to accurately assess confidence in their predictions while preserving their performance on relevant downstream tasks. Self-REF integrates naturally with any LLM backbone, allowing the model to condition its assessment of confidence on its generated outputs. It consists of three steps: (i) Confidence token annotation: the training dataset is generated by labeling each instance with confidence tokens based on the correctness of the base LLM. When the base LLM provides correct responses to the input text, we augment the instances with confident (<CN>) tokens. Conversely, if it responds incorrectly, we augment the instances with unconfident (<UN>) tokens. (ii) Self-REF fine-tuning: The base LLM is fine-tuned on the augmented training data, so that one would expect a confidence token to be generated following its predicted output. (iii) Confidence score extraction: To obtain a continuous confidence score, one computes the probability of the <CN> token, normalized by the sum of the probabilities of the <UN> and <CN> tokens.
@@ 2 Related Work
In this section, we briefly introduce related studies on uncertainty quantification, LLM routing, and LLM rejection learning. More discussion and details about related work are provided in Appendix A.
@@ 2.1 Uncertainty Quantification in LLMs
Recent studies in LLM uncertainty have explored leveraging internal features like logits and attention scores to estimate uncertainty without additional training (Zhou et al., 2023; Mahaut et al., 2024; Turpin et al., 2024). Post-processing methods such as temperature scaling (Guo et al., 2017), prompting LLMs to verbalize confidence scores (Tian et al., 2023), iterative prompting (Abbasi Yadkori et al., 2024), and instruction learning/pretraining (Zhang et al., 2021; Neeman et al., 2023; Li et al., 2023; Liu et al., 2024; Cohen et al., 2024) have also shown improvements in calibration. In contrast to these methods, Self-REF allows the LLM to learn to estimate its confidence at fine-tuning time. Furthermore, this internal notion of confidence is aligned against correctness rather than the consistency of responses after re-sampling, as is studied in (Yona et al., 2024; Manakul et al., 2023; Lin et al., 2023). As noted in (Huang et al., 2023; Wang & Zhou, 2024; Chuang et al., 2025), uncertainty scores derived from LLMs may exhibit a weak correlation with actual prediction correctness, and the most consistent answers are often not the most correct. Finally, while much of the prior work focuses on calibration, our work goes beyond it and focuses on the downstream utility of confidence scores for LLM routing and rejection learning.
@@ 2.2 LLM Routing
In confidence-based query-routing scenarios, one line of research trains additional classifiers to route queries to LLMs based on observed performance metrics and routing data, forming routing pipelines with trained multi-agent systems or extra routers to boost performance (Ding et al., 2024; Ong et al., 2024; Stripelis et al., 2024; Jiang et al., 2024; Zhang et al., 2023). Another approach leverages LLM cascades, invoking increasingly complex LLMs depending on trained scoring functions which predict whether an answer is correct, and consistency of responses upon resampling (Chen et al., 2023b; Yue et al., 2023). Unlike prior work in LLM cascades, Self-REF trains the estimate of the probability of correctness into the LLM itself instead of an external model. Here, the estimates of confidence from Self-REF are aligned against the correctness of the predictions, rather than the consistency. Additionally, Self-REF computes the confidence in the autoregressive decoding step, allowing for LLM itself to condition on its generated output. This dynamic approach allows for more accurate and adaptive routing decisions based on real-time confidence assessments.
@@ 2.3 LLM Rejection Learning
Learning to reject (Cortes et al., 2016) was originally introduced to equip the Bayes classification model with a rejection option for predictions. Despite advancements in LLMs, their predictions can still be prone to uncertainty due to inherent knowledge limitations. This highlights the need for LLMs to have the ability to refuse to answer, especially in situations where it is infeasible to fallback upon more advanced models. LLMs can be trained for the rejection task through rejection knowledge injection (Chen et al., 2024; Zhang et al., 2024), an additional rejection LLM agent (Mao et al., 2023, 2024), and rejection-oriented loss adaptations (Li et al., 2024; Mohri et al., 2024). However, such approaches involving extra knowledge injection or loss adaptation may degrade the base performance of the LLM. In this work, we aim to equip LLMs with the capability for rejection learning based on their confidence levels, without the need for designing new loss functions, thereby preventing any potential degradation in the performance of downstream tasks.
@@ 3.2 Self-REF
The Self-REF framework fine-tunes a base model $\mathcal{M}{}$ to use confidence tokens and also learn their embeddings. It involves (i) confidence token annotation, where data are augmented with confidence tokens, (ii) Self-REF fine-tuning on the augmented data, and (iii) confidence score extraction.
@@
0:  Base LLM $\mathcal{M}{}(\cdot)$ and original training data $\mathcal{D}_{\mathrm{train}}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N_{\text{train}}}$.
0:  Augmented data $\mathcal{D}^{\prime}_{\mathrm{train}}$ with unconfident samples and confident samples. Use $\mathcal{M}{}(\cdot)$ to make predictions $\widehat{y}^{(i)}$ for the given input query $x^{(i)}$.
1:  Create a set of unconfident samples:
$\mathcal{D}^{\prime}_{\mathrm{train},\texttt{<UN>}}=\{(x^{(i)},\widehat{y}^{(i)}\texttt{<UN>}):y^{(i)}\neq\widehat{y}^{(i)}\}_{i=1}^{N_{\text{train}}}.$
2:  Create a set of confident samples:
$\mathcal{D}^{\prime}_{\mathrm{train},\texttt{<CN>}}=\{(x^{(i)},\widehat{y}^{(i)}\texttt{<CN>}):y^{(i)}=\widehat{y}^{(i)}\}_{i=1}^{N_{\text{train}}}.$
3:  Mix the unconfident and confident data to form the combined augmented dataset, subsampling the unconfident samples with tunable proportion $\alpha\in[0,1]$:
@@
Note that the unconfident and confident tokens are generated end-to-end conditioned on both the prefix query and the answer that the LLM itself generated, in contrast to prior work (Ong et al., 2024; Ding et al., 2024; Stripelis et al., 2024) which typically routes based on the query alone with an extra router. By fine-tuning with the LLM’s own predictions on the augmented training set, Self-REF customizes the notion of uncertainty as the indicator for routing and rejection to the LLM itself, rather than relying on the inherent uncertainty of a query.
@@
After training, one can extract a continuous confidence score $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\in[0,1]$ by getting the token probability of <CN>, normalized over the sum of the <UN> and <CN> token probabilities: $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})=\frac{P(\texttt{<CN>})}{P(\texttt{<UN>})+P(\texttt{<CN>})}$. This continuous score gives us some control over navigating different tradeoffs in downstream settings.
@@ 3.3 Learning to Route and Reject
In confidence-based routing, answers where the smaller fine-tuned LLM is uncertain, i.e., $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$ for some chosen threshold $t\in[0,1]$, will be routed to a more powerful LLM $\mathcal{M}_{\text{Large}}(\cdot)$. For answers with $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\geq t$ (i.e., certain), the answers from the local LLM $\mathcal{M}{}(\cdot)$ will be directly returned to the user.
@@
To study this question, we create an evaluation set where half of the samples do not contain the ground truth, i.e., we remove all ground-truth information from $x^{(i)}$, and replace the label with “none of the above,” $y^{(i)}=\emptyset$. Ideally, the model’s confidence in any of the remaining answers should be low, and we can evaluate how good a policy for abstaining from answering the question is against this dataset. Thus, if $\mathcal{M}{}(\cdot)$ has low confidence, i.e., $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$ for a chosen threshold $t\in[0,1]$, we treat it as if the model were to abstain from answering (i.e., $\widehat{y}^{(i)}=\emptyset$).
@@ 4.1 Datasets and Baselines
Datasets
All experiments are conducted on the following four public datasets (more details in Appendix B):
  • MMLU (Hendrycks et al., 2021a, b): The Massive Multitask Language Understanding (MMLU) dataset consists of multiple-choice questions covering a wide range of knowledge domains with 57 distinct tasks.
  • OpenbookQA (Mihaylov et al., 2018): The OpenBookQA dataset is a question-answering dataset designed to test deep understanding through multi-step reasoning, common knowledge, and text comprehension, similar to open-book exams.
  • GSM8K (Cobbe et al., 2021): GSM8K is a dataset containing graduate school math questions, where each question is collected from the Math World Problem Repository (Roy & Roth, 2015).
  • MedQA (Jin et al., 2021): MedQA is a multiple-choice open domain question-answering dataset for problems collected from professional medical board exams.
@@
Figure 2: Overall accuracy vs. routing rate in confidence-based routing task. Routing is from local LLMs Llama3-8B-Instruct (row 1) and Mistral-7B-Instruct (row 2) to Llama3-70B-Instruct on MMLU (column 1), OpenbookQA (column 2), GSM8K (column 3), and MedQA (column 4). Routing rate of 0 gives pure local LLM performance, and routing rate of 1 gives pure Llama3-70B-Instruct performance. The solid black line corresponds to a baseline which chooses to route uniformly at random from the local fine-tuned LLM to the large LLM. Self-REF consistently outperforms baselines.
Baselines
We compare Self-REF to four state-of-the-art baselines for measuring model confidence, following the settings from existing work. Details about the prompts and hyper-parameters utilized in these baselines are given in Appendices B and C. The baselines are listed as follows:
  • Verbalizing Uncertainty (Tian et al., 2023): Prompt the LLM with in-context learning to yield confidence scores between 0 and 1. This encourages the LLM to generate its response while simultaneously estimating the confidence level of the response.
  • Verbalizing Yes/No Tokens (Tian et al., 2023): Prompt the LLM with a question of whether the model is confident in its answers. The confidence score is calculated by normalizing the probabilities of the “Yes” and “No” tokens by their sum.
  • Zero-shot Logits (Mahaut et al., 2024): Extract the probability of the generated prediction from the LLM without fine-tuning. In free-text generation questions, we average the probabilities of the tokens from the output answers. In the closed-form generation questions with single-token answers (e.g., multiple-choice questions), we directly select the probability of the predicted token.
  • Fine-tuned Logits (Liu et al., 2024): Extract the probability of the generated prediction from a fine-tuned LLM. All methods for computing the probabilities of the outputs are identical to the “Zero-shot Logits” baseline. The confidence score is determined by the probabilities.
Taking these continuous-valued confidence scores from each baseline, one can apply thresholds and compare the utility of these scores versus those of Self-REF in the downstream routing and rejection learning settings.
4.2 Experiment Settings
Confidence-based Routing Task
In the routing task, we aim to accurately route instances that the smaller LLMs (Llama3-8B-Instruct and Mistral-7B-Instruct) struggle with to the larger LLMs (Llama3-70B-Instruct) for improved performance. All experiments utilize Llama3-70B-Instruct with only its strong in-context learning capabilities during instance routing (no further fine-tuning). The routing decisions are determined by thresholding the normalized probability of the <CN> confidence token over the sum of the probabilities of <CN> and <UN> generated from Self-REF.
To characterize the trade-off between cost and performance, we analyze performance along several possible routing thresholds $t$. In particular, for a set of confidence scores $\mathcal{C}_{\mathcal{M},\mathcal{D}}=\{c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})\}_{i=1}^{N}$, we set $t$ at quantiles, where $t=\mathcal{Q}_{p}\left(\mathcal{C}_{\mathcal{M},\mathcal{D}}\right)$ for $p\in\{0,1,2,\dots,20\}$. Instances are routed to the larger LLM when $c_{\mathcal{M}}(x^{(i)},\widehat{y}^{(i)})<t$, so these thresholds correspond to routing approximately 0%, 5%, 10%, …, 100% of total queries to the larger LLM.